# On the Element-Wise Representation and Reasoning in Zero-Shot Image Recognition: A Systematic Survey

### v2

@@ III-A Problem Formulation
Given the seen domain $\mathcal{D}^{s}=\{(x^{s},y^{s},a^{s})|x^{s}\in\mathcal{X}^{s},y^{s}\in\mathcal{Y}^{s},a^{s}\in\mathcal{A}^{s}\}$, where $\mathcal{X}^{s}$, $\mathcal{Y}^{s}$, and $\mathcal{A}^{s}$ denote visual samples, category labels, and semantics (e.g., a set of attributes), and similarly, let $\mathcal{D}^{u}=\{(x^{u},y^{u},a^{u})|x^{u}\in\mathcal{X}^{u},y^{u}\in\mathcal{Y}^{u},a^{u}\in\mathcal{A}^{u}\}$ denote the unseen domain. Notably, $a^{s}\cup a^{u}\in\mathbb{R}^{m\times k}$ are class-level semantics annotations, where $m$ denotes the number of categories and $k$ is the number of attributes. Without loss of generality, the task of ZSOR can be modeled as learning a mapping/relational function $\Psi:\mathcal{X}\to\mathcal{A}$, wherein $\mathcal{X}^{u}$ is strictly inaccessible for training. During inference, the learned function $\Psi$ is applied to recognize samples from the unseen domain only, i.e., ZSL, or from the joint of both seen and unseen domains, i.e., Generalized ZSL (GZSL).
@@
The success of ZSOR relies on the sharing property between $\mathcal{A}^{s}$ and $\mathcal{A}^{u}$, which act as the bridge from seen to unseen domains. An intuitive idea is to learn the visual representations corresponding to attributes, which is the notion of element-wise learning. However, due to the single target and relatively closed environment of object recognition, learning for $\Psi$ can be transformed into a plain class-wise distribution matching problem. Class-wise distribution matching concerns neither the discrepancies of the bottom attributes nor the variations of fine-grained vision. It treats both visual and attribute features as a whole, and the objective is to reduce the distributional distance between the two so as to learn an unbiased mapping function. Mainstream techniques include semantic space design, generative methods and regularization. The goal of semantic space design is to search for a suitable low-dimensional space that perfectly matches visual and attribute embeddings. Common practices are one-way mapping [50, 51], which projects visual/attribute into the space of the other, and common space [44], which projects both visual and attribute into a common space. The goal of the generative approach is to simulate the visual features of unseen classes [52, 53]. The basic approach is to train a generator that produces visual features based on attributes with the help of networks such as GANs [54] and VAEs [55]. The goal of regularization is imposing constraints to improve the mapping performance or robustness of the model. A popular approach is to design a specific loss function. For example, Liu et al. [56] uses an information entropy loss for the unseen class thereby reducing uncertainty.
Despite the impressive results achieved by these methods, they provide limited insights for micro-operations and bottom-up analyses, and are therefore out of the scope of this paper. We recommend the reader to refer to other surveys [24, 6] for a deeper understanding of the above methods.
@@
Depending on who the method operates on, we categorize the element-wise techniques into three broad classes, including Visual Component Analysis, Visual-Attribute Matching, and Attribute Component Analysis. Then the methods are further categorized into different subclasses based on the specific techniques they take.
III-B Visual Component Analysis
@@
Visual component analysis is the refined dissection of an image or its features, involving localization, modeling, and so on. Images often contain a large amount of redundant information, only a few of which are decisive for discrimination. At the same time, images lack fine-grained annotations. Therefore, the goal of visual component analysis is to reduce the interference of irrelevant information as well as noise, and prompt the model to focus on those critical visual parts. Commonly used methods can be classified into the following three categories: local visual attention, data manipulation and graph modeling.
III-B1 Local Visual Attention
@@
People selectively focus on the most attractive parts of things when they look at them, and these places usually contain the most discriminative information. It is this behavior that the attention mechanism mimics to make the model concentrate on the most significant areas, thus improving the efficiency and accuracy of information processing. Attention mechanisms have a wide range of applications in computer vision, natural language processing and other fields. The goal of local visual attention is to utilize the attention mechanism to identify the most valuable visual regions and reduce the influence of background and irrelevant regions. Specifically, local visual attention generates multiple masks for visual features. Each mask is a normalized weight matrix that reveals a significant region. Enhanced features can be obtained by weighting the original features using mask, thus improving the discriminability (Fig. 3). Suppose that $x\in\mathbb{R}^{C\times H\times W}$ denotes the visual feature, where $C$, $H$, and $W$ are the channel, height, and weight, respectively. Let $M=\mathcal{F}(x)$ represents the attention mask, where $\mathcal{F}$ denotes a small learnable network, e.g., $1\times 1$ convolutional network, and $M\in\mathbb{R}^{N\times H\times W}$ denotes $N$ attention masks. Afterward, multiplying the masks with the original features yields $N$ regional features, which can be expressed as:
@@
III-B2 Data Manipulation
The objective of data manipulation is the same as local visual attention, i.e., to search and localize the most critical regions in an image. The difference is that local visual attention is weighted at the semantic feature level, whereas data manipulation operates at the image level. An intuitive scheme is to recognize parts of an object, such as the head, body and claws of a bird, with the help of detection techniques. The detection network can provide coordinates of multiple parts simultaneously. Afterwards, image patches are cropped based on the coordinates or the relevant locations are localized in deep features. These parts introduce a large amount of fine-grained semantic information that assists in facilitating the classification ability of the model  [58, 59]. Additionally, the technique also improves the performance of generative methods, which struggle to capture fine-grained visual information differences. For example, AGAA [60] utilizes a detector to extract region information that is then sent to a discriminator for training along with class embeddings generated by a generator. The application of the detection technique is very limited because the performance of the detector relies on part-level visual annotations, which is undesirable in most scenarios.
@@ III-B2 Data Manipulation
where $\mathcal{F}$ is a learnable network. $Z$ is the window, $z_{h},z_{w}$ denote the coordinates, and $z_{l}$ denotes the length of the region. After that, the regions located at the coordinates are zoomed in and re-fed to the classification network. SGMA [35], on the other hand, unites local visual attention techniques to generate multiple region coordinates at once. It first generates multiple attention masks and then uses the attention masks as inputs to get the corresponding coordinates. Finally, the original image is cropped according to the coordinates. Furthermore, SR2E [62] and ERPCNet [63] introduce the reinforcement learning paradigm, which induce the cropping behavior of the model by setting up special rewards and penalties to obtain the most discriminative sub-image. POPRNet [64] takes unsupervised detection technique. It generates multiple proposal regions for an image and then sorts and selects Top-K region features based on category confidence. ZSLViT [65] utilizes the patch split mechanism of ViT [66] to get multiple region features and later filters the key visual regions by calculating the attention weights between the regional and class tokens.
Unlike the above approaches, to capture the key visual semantic embeddings, VGSE-SMO [67] takes a patch clustering approach. It utilizes an unsupervised compact watershed segmentation algorithm [68] to decompose images into multiple patches, after which individual patches are linked to class labels via a classification network. By doing so, the model implicitly learns the relationship between classes and fine-grained visuals.
III-B3 Graph Modeling
Visual region localization promotes a model that focuses on important local features, however, fails to take into account potential relationships between regions. For example, a bird has biological topological connections between its various body parts. Independent regional representation loses substantial semantic information by narrowing the field of view, which may lead to semantic ambiguity or disambiguation. Sharing information between regions can be a useful method to correct semantic information by compensating for the loss of field of view. Fortunately, graph modeling techniques is able to well handle such irregular non-geometric data with a structured message propagation mechanism [7]. In recent years, with the birth and development of graph convolutional networks (GCNs) [69], graph modeling techniques have been widely used in various computer vision tasks. Graph convolutional networks ingeniously integrate the concept of graph into convolutional neural networks. Suppose $W^{(l)}$ denotes the parameters of the $l$-th layer of GCNs, the output of the $(l+1)$-layer can be expressed as:
$H^{(l+1)}=\sigma(\tilde{D}^{-\frac{1}{2}}\tilde{A}\tilde{D}^{-\frac{1}{2}}H^{(l)}W^{(l)}),$ (3)
where $A$ is the adjacent matrix and $\tilde{D}$ denotes its degree matrix. $\sigma$ denotes the activation function, and $H^{(l)}$ is the output of the $l$-layer of GCNs. It can be seen that the nodes share information through an inside-out information aggregation mechanism through the graph network.
@@ III-B3 Graph Modeling
III-C Visual-Attribute Matching
The objective of visual-attribute matching is to find appropriate visual features for each attribute, and to learn fine-grained element-wise representations by searching and matching pairs of visual-attribute embeddings. Compared to visual component analysis, visual-attribute matching has a more explicit motivation to interact with attributes. At the same time, visual-attribute matching techniques have a stronger perception of subtle variations between attributes, leading to better generalization to unseen domains. Such text-induced visual element learning is more compatible with human learning habits and more challenging. Existing approaches can be categorized into cross attention and prototype learning.
III-C1 Cross Attention
@@
Attention mechanisms have been used in visual component analysis to localize discriminative regions that are friendly to class labels but fail to capture subtle visual differences between attributes. Therefore, cross attention is introduced to discover visual regions of interest for attributes. Cross attention allows for simultaneous processing of different sources of information and efficient modeling of their dependencies. Among them, dot product attention, due to its high computational efficiency and low space cost is widely applied in many advanced models such as Transformer [72]. The core of dot product attention is to compute the similarity between query and key-value pairs to assign attention weights, and the similarity measure is the dot product of vectors. Given a query series $Q$, a key series $K$ and a value series $V$, the attention score can be formulated as:
$Attention(Q,K,V)=softmax(\frac{QK^{\mathsf{T}}}{\sqrt{d_{k}}})V,$ (4)
where $d_{k}$ denotes the dimension of the key vector which is used to scale the dot product result. When $Q$, $K$, and $V$ all come from the same input series, the attention score learns the dependencies between the parts of the series, i.e., self-attention. And when there are two different input series involved in the computation, the attention score learns the dependencies between the different series, i.e., cross attention.
In order to uncover the interactions between attributes and visual regions, the cross attention technique is naturally introduced into ZSOR (Fig. 4). For example, DAZLE [39] proposes a dense attention technique that assigns weights to each region based on attributes. Specifically, it utilizes the features output from the last layer of the convolutional network and treats each point on the feature as a region. The similarity between attributes and regions is then computed with cross attention. Suppose $x\in\mathbb{R}^{C\times r}$ denotes the visual feature with $r=H\times W$ regions. Let $a\in\mathbb{R}^{d\times m}$ denote the attribute vector, where $m$ is the number of attributes and $d$ is the vector dimension. The cross attention score is computed as:
where $w\in\mathbb{R}^{d\times C}$ is a learnable matrix which ensures the visual and attribute vectors are in the same dimension space. The weighted visual features reflect attribute preferences. After that DAZLE uses the attention scores of all the attributes to classification. GEM [73] employs the same approach, however it retains the maximum attention scores with max-pooling during classification. Composer [74] synthesizes the features of unseen classes with critical regions identified by attention mechanism, while AREES [75] exploits the regional features to improve the performance of VAE.
Some studies resort to ViT [66] to model dependencies between vision and attributes [76, 77, 78]. The essential cross attention mechanism is identical. The difference is that ViT divides the image into patches, each of which is considered as a visual region.
@@
In addition, bidirectional attention is explored in depth. It is important to know not only which region the attribute is attending to, but also which attribute the region is attending to. Based on this idea, several methods are developed. For example, MSDN [79] proposes a mutual distillation network to align attribute-region attention. The attribute-to-region attention algorithm is the same as Eq. 5, and the region-to-attention to attributes is formulated as:
$Attention(x,a)=\frac{\exp(x^{\mathsf{T}}wa)}{\sum^{m}\exp(x^{\mathsf{T}}wa)}a,$ (6)
where $w\in\mathbb{R}^{C\times d}$ is a learnable matrix. Afterwards, it uses two attention scores to compute the respective classification results and aligns them with distillation loss. Ge et al. [48] also integrate bidirectional attention techniques to improve the visual-attribute matching ability of model.
III-C2 Prototype Learning
@@
The objective of prototype learning is to store representative samples and perform tasks such as classification and regression by calculating distances to them [80]. A prototype is a standardized sample or example that represents a category or type. For example, the mean vector of a collection of vectors or the center of mass of a distribution is a prototype. Prototype learning imitates the human mindset of analogical reasoning, i.e., reasoning and summarizing by analogy with typical examples. Its main advantage is its stability in dealing with complex data distributions, especially when there is overlap or imbalance between data categories. Therefore, for tasks such as out-of-distribution detection [81] and low-shot learning [82], prototype learning is considered to have stronger robustness.
@@
For visual-attribute matching, there is semantic overlap in the visual region. The same region may contribute to multiple attributes. Although cross attention is able to measure the contribution value by similarity, the lack of an explicit means of visual decoupling may lead to pseudo-correlation of visual-attributes. Therefore, learning a separate visual prototype for each attribute is a promising solution. When the prototype is acquired, the visual region can decide which attribute it belongs to by calculating the distance between the region vector and the prototype. It can be observed that cross attention discovers pairing relations through the similarity between attribute embeddings and visual embeddings, while prototype learning discovers visual-attribute pairing relations indirectly through the similarity between visual embeddings and visual prototype embeddings.
@@
APN [83] first introduces prototype learning to visual-attribute matching. It initializes a learnable prototype vector for each attribute and then calculates the similarity between the visual region and the prototype. After that, the largest region vector is taken as the attribute vector, and then regresses with the attribute embedding. In addition, it incorporates a decorrelation loss [84] to ensure the independence between individual attributes. DPPN [85] considers that the prototype possesses a seen-domain preference, which cannot capture the delicate instance-level attribute shift. Therefore, it is necessary to equip each instance with exclusive attribute prototypes. It proposes a progressive attribute localization method, which dynamically adjusts the prototype by locating attributes through the prototype while updating the prototype in turn. In addition,  [48, 86, 87, 88] also achieve favorable results by using prototype learning.
III-D Attribute Component Analysis

### v3

@@ 3.1 Problem Formulation
Given the seen domain $\mathcal{D}^{s}=\{(x^{s},y^{s},a^{s})|x^{s}\in\mathcal{X}^{s},y^{s}\in\mathcal{Y}^{s},a^{s}\in\mathcal{A}^{s}\}$, where $\mathcal{X}^{s}$, $\mathcal{Y}^{s}$, and $\mathcal{A}^{s}$ denote visual samples, category labels, and semantics (e.g., a set of attributes), and similarly, let $\mathcal{D}^{u}=\{(x^{u},y^{u},a^{u})|x^{u}\in\mathcal{X}^{u},y^{u}\in\mathcal{Y}^{u},a^{u}\in\mathcal{A}^{u}\}$ denote the unseen domain. Notably, $a^{s}\cup a^{u}\in\mathbb{R}^{m\times k}$ are class-level semantics annotations, where $m$ denotes the number of categories and $k$ is the number of attributes. Without loss of generality, the task of ZSOR can be modeled as learning a mapping/relational function $\Psi:\mathcal{X}\to\mathcal{A}$, wherein $\mathcal{X}^{u}$ is strictly inaccessible for training. During inference, the learned function $\Psi$ is applied to recognize samples from the unseen domain only, i.e., ZSL, or from both seen and unseen domains, i.e., Generalized ZSL (GZSL).
@@
The success of ZSOR relies on the shared attribute between $\mathcal{A}^{s}$ and $\mathcal{A}^{u}$, which act as the bridge between seen and unseen domains. The intuitive idea is to learn the visual representations corresponding to attributes, which is the notion of element-wise learning. However, due to the single target and relatively closed environment of object recognition, learning for $\Psi$ can be transformed into a plain class-wise distribution matching problem. Class-wise distribution matching concerns neither the discrepancies of the bottom attributes nor the variations of fine-grained vision. It treats both visual and attribute features as a whole, and the objective is to reduce the distributional distance between the two to learn an unbiased mapping function. Mainstream techniques include semantic space design, generative approach, and regularization. Semantic space design is to search for a suitable low-dimensional space that perfectly matches visual and attribute embeddings [49, 43]. The generative approach is to simulate the visual features of unseen classes [50, 51] by training a generator that produces visual features based on attributes with the help of networks such as GANs [52, 53, 54], VAEs [55, 56] and generative flows [57, 58]. Regularization is imposing constraints to improve the mapping performance or robustness of the model. For example, Liu et al. [59] use an information entropy loss to reduce uncertainty. Despite the impressive results achieved by these methods, they provide limited insights for micro-operations and bottom-up analyses and are, therefore, out of the scope of this survey. The readers can refer to other related surveys [23, 24] for a deeper understanding.
@@
Depending on the concrete elements the method operates on, we categorize the element-wise techniques into three broad categories, i.e., Visual Component Analysis, Visual-Attribute Matching, and Attribute Component Analysis. Then, the methods are further categorized into different sub-categories based on the specific techniques they use.
3.2 Visual Component Analysis
@@
Visual component analysis is the refined dissection of an image or its features, involving localization, modeling, and so on. Images often contain a large amount of redundant information, only a few of which are decisive for discrimination. At the same time, images lack fine-grained annotations. Therefore, the goal of visual component analysis is to reduce the interference of irrelevant information as well as noise, and prompt the model to focus on those critical visual parts. Commonly used methods can be classified into the following three sub-categories: Local Visual Attention, Data Manipulation, and Graph Modeling.
3.2.1 Local Visual Attention
@@
Attention mechanisms are effective to improve efficiency and accuracy in many visual tasks by mimicking human focuses. Motivated by this, local visual attention is to utilize the attention mechanism to identify the most valuable visual regions and reduce the influence of background or irrelevant regions. Specifically, local visual attention generates multiple masks for visual features. Each mask is a normalized weight matrix that reveals a significant region. Enhanced features can be obtained by weighting the original features using mask, thus improving the discriminability (Fig. 3). Suppose that $x\in\mathbb{R}^{C\times H\times W}$ denotes the visual feature, where $C$, $H$, and $W$ are the channel, height, and weight, respectively. Let $M=\mathcal{F}(x)$ represents the attention mask, where $\mathcal{F}$ denotes a small learnable network, e.g., $1\times 1$ convolutional network, and $M\in\mathbb{R}^{N\times H\times W}$ denotes $N$ attention masks. Afterward, multiplying the masks with the original features yields $N$ regional features, which can be expressed as:
@@
3.2.2 Data Manipulation
Data manipulation has a same objective as local visual attention, i.e., to search and localize the most critical regions. The difference is that the latter works at the feature space, whereas the former works at the image level. An intuitive approach is employing detection techniques to recognize parts of an object, such as a bird’s head, body, and claws. The detection network can provide coordinates for multiple parts simultaneously, allowing image patches to be cropped accordingly or deep features to be localized. This method introduces fine-grained semantic information that enhances the model’s classification ability [61, 62]. Additionally, it boosts the performance of generative methods, which often struggle with capturing fine-grained visual differences. For example, AGAA [63] utilizes a detector to extract region information that is fed into a discriminator, along with class embeddings generated by a generator. However, this technique is limited in practice since the detector’s performance heavily depends on part-level visual annotations, which are often impractical to obtain.
@@ 3.2.2 Data Manipulation
where $\mathcal{F}$ is a learnable network. $Z$ is the window, $z_{h},z_{w}$ denote the coordinates, and $z_{l}$ denotes the length of the region. After that, the regions located at the coordinates are zoomed in and re-fed to the classification network. SGMA [34] incorporates local visual attention techniques to coordinates generation. Furthermore, SR2E [65] and ERPCNet [66] introduce the reinforcement learning paradigm, which induce the cropping behavior of the model by setting up special rewards and penalties to obtain the most discriminative sub-image. POPRNet [67] takes unsupervised detection technique. It generates multiple proposal regions for an image and then sorts and selects Top-K region features based on category confidence. ZSLViT [68] utilizes the patch split mechanism of ViT [69] to get multiple region features and later filters the key visual regions by calculating the attention weights between the regional and class tokens.
Unlike the above approaches, to capture the key visual semantic embeddings, VGSE-SMO [70] takes a patch clustering approach. It utilizes an unsupervised segmentation algorithm to decompose images into multiple patches, after which individual patches are linked to class labels via a classification network. By doing so, the model can implicitly learn the relationship between classes and fine-grained visual patches that align better with each other.
3.2.3 Graph Modeling
Visual region localization helps models focus on important local features but often overlooks the potential relationships between regions. For example, a bird’s body parts are biologically connected, and treating them independently can cause semantic ambiguity due to a narrowed field of view. Sharing information between regions can mitigate this by preserving essential semantic context. Graph modeling techniques are well suited for handling such irregular non-geometric data with a structured message propagation mechanism [17]. In recent years, with the birth and development of graph convolutional networks (GCNs) [71], graph modeling techniques have been widely used in various computer vision tasks.
@@ 3.2.3 Graph Modeling
3.3 Visual-Attribute Matching
The goal of visual-attribute matching is to align visual features with their corresponding attributes by learning fine-grained, element-wise representations through the pairing of visual-attribute embeddings. Unlike visual component analysis, this approach explicitly focuses on interacting with attributes, allowing for a better perception of subtle attribute variations and improved generalization to unseen domains. This text-driven visual learning aligns more closely with human learning patterns and presents a greater challenge. Current methods in this area can be categorized into two sub-categories: Cross Attention and Prototype Learning.
3.3.1 Cross Attention
Cross attention allows for the simultaneous processing of different sources of information and efficient modeling of their dependencies. Among them, dot product attention, due to its high computational efficiency and low space cost, is widely applied in many advanced models such as Transformer [74]. The core of dot product attention is to compute the similarity between query and key-value pairs to assign attention weights, and the similarity measure is the dot product of vectors. Given a query series $Q$, a key series $K$ and a value series $V$, the attention score can be formulated as:
$Attention(Q,K,V)=softmax(\frac{QK^{\mathsf{T}}}{\sqrt{d_{k}}})V,$ (3)
where $d_{k}$ denotes the dimension of the key vector used to scale the dot product result. When $Q$, $K$, and $V$ all come from the same input series, the attention score learns the dependencies between the parts of the series, i.e., self-attention. When there are two different input series involved in the computation, the attention score learns the dependencies between the different series, i.e., cross attention.
To uncover the interactions between attributes and visual regions, the cross attention technique is naturally introduced into ZSOR (Fig. 4). For example, DAZLE [38] proposes a dense attention technique that assigns weights to each region based on attributes. Specifically, it utilizes the features output from the last layer of the convolutional network and treats each point on the feature as a region. The similarity between attributes and regions is then computed with cross attention. Suppose $x\in\mathbb{R}^{C\times r}$ denotes the visual feature with $r=H\times W$ regions. Let $a\in\mathbb{R}^{d\times m}$ denote the attribute vector, where $m$ is the number of attributes, and $d$ is the vector dimension. The cross attention score is calculated as:
where $w\in\mathbb{R}^{d\times C}$ is a learnable matrix that ensures the visual and attribute vectors are in the same dimension space. The weighted visual features reflect attribute preferences. After that, the attention scores of all the attributes are used for classification. GEM [75] employs the same approach, however it retains the maximum attention scores with max-pooling during classification. Composer [76] synthesizes the features of unseen classes with critical regions identified by attention mechanism, while AREES [77] exploits the regional features to improve the performance.
@@ 3.3.1 Cross Attention
Some studies resort to ViT [69] to model dependencies between vision and attributes [78, 79, 80]. The essential cross attention mechanism is identical. The difference is that ViT divides the image into patches, each of which is considered a visual region. In addition, bidirectional attention is explored in depth. It is important to know not only which region the attribute is attending to, but also which attribute the region is attending to. Based on this idea, several methods are developed. For example, MSDN [81] proposes a mutual distillation network to align attribute-region attention. Ge et al. [47] also integrate bidirectional attention techniques to improve the visual-attribute matching ability of the model.
3.3.2 Prototype Learning
@@
The gist of prototype learning is to perform tasks like classification and regression using a set of representative samples, known as prototypes [82] to regularize the learning process. A prototype is a standardized example that captures the essence of a category or type, such as the mean vector of a set of vectors or the center of mass of a distribution. Prototype learning mimics human analogical reasoning by drawing conclusions based on typical examples. Its main advantage is its robustness in handling complex data distributions, particularly when there is overlap or imbalance between data categories. As a result, prototype learning is considered to be more effective for tasks like out-of-distribution detection [83] and low-shot learning [84].
@@
For visual-attribute matching, there is semantic overlap in the visual region. The same region may contribute to multiple attributes. Although cross attention is able to measure the contribution value by similarity, the lack of an explicit means of visual decoupling may lead to pseudo-correlation of visual-attributes. Therefore, learning a separate visual prototype for each attribute is a promising solution. When the prototype is acquired, the visual region can decide which attribute it belongs to by calculating the distance between the region vector and the prototype. It can be observed that cross attention discovers pairing relations through the similarity between the attribute and visual embeddings, while prototype learning discovers visual-attribute pairing relations indirectly through the similarity between the embeddings of samples and prototypes.
@@
APN [85] first introduces prototype learning to visual-attribute matching. It initializes a learnable prototype vector for each attribute and then calculates the similarity between the visual region and the prototype. After that, the largest region vector is taken as the attribute vector, and then regresses with the attribute embedding. In addition, it incorporates a decorrelation loss [86] to ensure the independence between individual attributes. DPPN [87] considers that the prototype possesses a seen-domain preference, which cannot capture the delicate instance-level attribute shift. Therefore, it is necessary to equip each instance with exclusive attribute prototypes. It proposes a progressive attribute localization method, which dynamically adjusts the prototype by locating attributes through the prototype while updating the prototype in turn. In addition, some other methods [47, 88, 89, 90, 91] also achieve favorable results by using prototype learning.
3.4 Attribute Component Analysis